# Nonuniqueness of capped black holes: large and small bubbles

### v1

@@ IV Physical properties
Finally, we study the phase space of the new solution. The new capped black hole solution is determined by four independent parameters $(\ell,\sigma,\tilde{\nu},\tilde{\gamma})$. Here, we focus on the phase diagram in the $\sigma=0.8$ case. Other cases are qualitatively the same.
@@ IV.1 Thermodynamics
$\Phi_{\mathcal{D}}$ $:=-i_{v_{\rm in}}A\biggr{|}_{x=1}=-(A_{\phi}+A_{\psi})\biggr{|}_{x=1}=-\frac{\sqrt{3}\ell cs\tilde{v}_{0}\left(2c(1-\gamma)f_{2}\nu\tilde{g}_{1}(a-b)-f_{1}s\left(c_{3}f_{4}f_{17}-bc_{2}f_{5}f_{13}\right)\right)}{2\nu(1-\gamma)f_{1}f_{2}},$ (92)
@@
with $\Phi(y):=-i_{v_{\cal H}}A|_{x=1}=-(A_{t}+\omega_{\psi}A_{\psi}+\omega_{\phi}A_{\phi})|_{x=1}$. Henceforth, the thermodynamics implies that the capped black hole is parametrized, at most, by the five independent parameters $(M,J_{\psi},J_{\phi},Q,\Phi_{\mathcal{D}})$. However, it is not clear whether these five parameters uniquely determine the solution since the known uniqueness theorem assumes more parameters to specify the solution Armas:2014gga .
Figure 4: The $\sigma=0.8$ phase shown by the colored regions in the $(j_{\psi},j_{\phi})$, $(\phi_{\mathcal{D}},j_{\phi})$ and $(q_{\mathcal{D}},j_{\phi})$ planes. The $\tilde{\nu}={\rm constant}$ phases are shown by thin blue curves with values of $\tilde{\nu}$. To keep the visibility in the panel (a), the $\tilde{\nu}={\rm constant}$ curves are displayed only in the close-up. In each panel, the extremal capped black hole limit, extremal black hole limit and $\tilde{\gamma}=0$ limit are shown by black, orange and black dashed curves, respectively. The red dot-dashed curves denote critical curves for large and small bubble branches, which also become a part of the boundary in the panel (a). In the panel (a), two branches exists for given $(j_{\psi},j_{\phi})$ in the purple region, while only the large bubble branch in the gray region. In the panel (b) and (c), the critical curves separate two branches and the orange dashed curves correspond to the borders between the single and double branch regions. In the panel (b) and (c), the red and blue regions represent to the small bubble branch and corresponding large bubble branch that overlaps in the purple region of the panel (a), respectively, and only the large bubble branch exists in the gray region. In each panel, the extremal capped black hole limit and extremal black hole limit intersect at $A$ and the $\tilde{\gamma}=0$ limit intersects with the former two at $B$ and $C^{\prime}$, respectively. The critical curve has an end on the extremal black hole limit at $E$ and another on the $\tilde{\gamma}=0$ limit at $D$. In the panel (c), the extremal capped black hole limit and $\tilde{\gamma}=0$ limit do not intersect, but instead, are connected by the $\partial_{\tilde{\nu}}j_{\phi}\partial_{\tilde{\gamma}}q_{\mathcal{D}}-\partial_{\tilde{\nu}}q_{\mathcal{D}}\partial_{\tilde{\gamma}}j_{\phi}=0$ curve in the $(q_{\mathcal{D}},j_{\phi})$ plane (blue dot-dashed curve), which intersect with the extremal capped black hole limit at $F$ and meet with the $\tilde{\gamma}=0$ at $C^{\prime}$.
@@
In Fig. 4, the phase for $\sigma=0.8$ is depicted in the $(j_{\psi},j_{\phi})$, $(\phi_{\mathcal{D}},j_{\phi})$, and $(q_{\mathcal{D}},j_{\phi})$ planes, where the solution exists within the colored regions. Since the BPS bound is saturated only at the limit $\sigma\to\pm 1$, all the phases for $\sigma=0.8$ are non-supersymmetric. In each panel, curves of constant $\tilde{\nu}$ are shown as thin blue lines. The capped black hole solutions are not uniquely determined by their conserved charges $(M,Q,J_{\psi},J_{\phi})$ and feature two distinct branches, which we call large bubble branch and small bubble branch. As explained in detail below, the former always has the larger bubble than the latter. In Fig. 4(a), both branches coexist within the purple region. These branches are distinguishable by their magnetic flux or potential values. In Figs. 4(b) and (c), the small bubble branch is represented by the red region, while the large bubble branch occupies both blue and gray regions. The portion of the large bubble branch that shares the same $(M,Q,J_{\psi},J_{\phi})$ as the small bubble branch is highlighted in blue. The segment that does not overlap with the small bubble branch in terms of $(M,Q,J_{\psi},J_{\phi})$ is shown in gray.
@@
  • Nonsupersymetric extremal black hole limit (orange solid curves in Fig. 4), as $\tilde{\gamma}\to\sqrt{2-\tilde{\nu}^{2}}$ with $\tilde{\nu}$ held constant, where the surface gravity $\kappa$ approaches zero. In this limit, the angular momenta correspond to those of the extremal Cvetič-Youm black holes, maintaining the same charge-to-mass ratio as specified in Eq. (88)
@@
    $=\frac{\left(\sigma^{3}+1\right)\left(3\left(\sigma^{3}-3\right)\tilde{\nu}^{2}(\tilde{\nu}^{2}-2)+2\left(\sqrt{\sigma^{6}\left(-3\tilde{\nu}^{4}+6\tilde{\nu}^{2}+1\right)+2\sigma^{3}\left(3\tilde{\nu}^{4}-6\tilde{\nu}^{2}+1\right)+1}-\sigma^{3}-1\right)\right)}{9\tilde{\nu}^{2}\left(2-\tilde{\nu}^{2}\right)}.$ (96)
    The entropy, magnetic flux and potential have the limit
@@
    ×$\quad\times\sqrt{\sigma^{6}\left(-3\tilde{\nu}^{4}+6\tilde{\nu}^{2}+1\right)+2\sigma^{3}\left(3\tilde{\nu}^{4}-6\tilde{\nu}^{2}+1\right)+1}\biggr{]}^{1/2},$ (97)
@@
    with $(j_{1},j_{2})$ in Eq. (96). Since the Cvetič-Youm black hole does not possess the bubble of $S^{2}$ outside the horizon, both the magnetic flux and magnetic potential are trivially zero. We note that the limit of the entropy (97) does not coincide with that of of the corresponding Cvetič-Youm black hole in Eq. (99). The similar discontinuity is reported for the black hole limit in the black ring phase Elvang:2007hs ; Elvang:2004rt . As speculated in Ref. Elvang:2004rt , this may be because the spacetime topology changes at the limit.
    The extremal capped black hole has two endpoints at $\tilde{\nu}\to 1$ and $\tilde{\gamma}\to 1$ ($A$ in Fig. 4) and $\tilde{\nu}\to\sqrt{2}$ and $\tilde{\gamma}\to 0$ ($B$ in Fig. 4), which have the dimensionless angular momenta and entropy, respectively,
    $(j_{\psi},j_{\phi})\bigr{|}_{A}$ $=\frac{(j_{1}+\sigma^{3}j_{2},j_{2}+\sigma^{3}j_{1})}{(1+\sigma^{2})^{3/2}}\biggr{|}_{A},$ (100)
    $j_{1}\bigr{|}_{A}$ $=1+j_{2}\bigr{|}_{A}=\biggr{\{}\begin{array}[]{cc}1-9^{-1}(2-\sigma^{3})^{2}&(2^{-1/3}<\sigma<1)\\ 1-\sigma^{6}&(-1<\sigma<2^{-1/3})\end{array},$ (103)
    $a_{{\cal H}}\bigr{|}_{A}$ $=2\sqrt{2}\left(\frac{1+\sigma^{3}}{1+\sigma^{2}}\right)^{3/2}\times\quad\biggr{\{}\begin{array}[]{cc}9^{-1}(2-\sigma^{3})\sqrt{2\sigma^{3}-1}&(2^{-1/3}<\sigma<1)\\ \sqrt{\sigma^{6}(1-2\sigma^{3})}&(-1<\sigma<2^{-1/3})\end{array}.$ (106)
@@
    $(j_{\psi},j_{\phi})\bigr{|}_{B}$ $=\frac{(1,\sigma^{3})}{(1+\sigma^{2})^{3/2}},$ (107)
    $a_{{\cal H}}\bigr{|}_{B}$ $=0.$ (108)
    In particular, the endpoint $B$ corresponds to the extremal Cvetič-Youm phase with zero entropy.
  • Nonsupersymetric, extremal capped black hole limit (black curves in Fig. 4) as $\tilde{\nu}\to 1$ and $\tilde{\gamma}\to 1$ with
    $\tilde{\gamma}=1-(\delta^{2}/\mu^{2})\beta,\quad\tilde{\nu}=1+\delta^{2}\beta,\quad(0<\mu<1).$ (109)
    $(\sigma,\tilde{b},\delta,\mu)$ which can be solved for a given set of $(\sigma,\mu)$. Substituting these expressions into Eqs. (84) and (87), and subsequently taking the limit $\beta\to 0$, we derive these two equations for the parameters $(\sigma,\tilde{b},\delta,\mu)$, solvable for a given $(\sigma,\mu)$ configuration. This phase intersects with the previous extremal black hole limit (orange solid curves) at $\mu\to 1$ ($A$ in Fig. 4) and terminates at $\mu\to 0$ ($C$ in Fig. 4). At the latter point, the spacetime corresponds to the extremal phase with zero entropy. For instance, at $\sigma=0.8$, the physical parameters are numerically determined as follows:
@@
    where $j_{\psi}$ has a local minimum on a $j_{\phi}={\rm constant}$ surface. This curve forms part of the boundary in the phase diagram only in the $(j_{\psi},j_{\phi})$ plane, as shown in Fig. 4(a). In the $(\phi_{\mathcal{D}},j_{\phi})$ and $(q_{\mathcal{D}},j_{\phi})$ planes, it distinguishes between the large and small bubble branches, as illustrated Fig. 4(b) and (c). This curve ends on the curve $\tilde{\gamma}=0$ ($D$ in Fig. 4) and on the curve of the extremal black hole limit ($E$ in Fig. 4). The respective physical quantities at these end points are numerically determined as
    $(j_{\psi},j_{\phi},\phi_{{\mathcal{D}}},q_{{\mathcal{D}}})\bigr{|}_{D}$ $=(0.461...,0.252...,0.817...,-0.00744...),$ (112)
    $(j_{\psi},j_{\phi},\phi_{{\mathcal{D}}},q_{{\mathcal{D}}})\bigr{|}_{E}$ $=(0.420...,0.187...,0.866...,0).$ (113)
    The $\tilde{\gamma}=0$ ($\gamma=1)$ limit (black dashed curves in Fig. 4). This is a nonextremal and nonsigular phase. This allows the potential possibility to extend the phase diagram beyond this curve, as discussed in Ref. Suzuki:2024vzq for the charge black ring. In such an extended region, $\gamma$ might exceed $1$, leading $(\beta,a,\tilde{b})$ to assume complex values, which we do not discuss here. This curve initiates at $\tilde{\nu}=\sqrt{2}$ at point $B$ in Fig. 4 and concludes at $\tilde{\nu}=1$ at point $C^{\prime}$ in Fig 4(c), where the parameters are given by:
    $(j_{\psi},j_{\phi},\phi_{{\mathcal{D}}},q_{{\mathcal{D}}})\bigr{|}_{C^{\prime}}=(0.482\ldots,0.240\ldots,0.789\ldots,-0.0430\ldots).$ (114)
@@
    where $q_{\mathcal{D}}$ has a local maximum on the $j_{\phi}={\rm constant}$ phase (Fig. 5). As in Fig. 4(c), this curve forms a part of the boundary in the $(q_{\mathcal{D}},j_{\phi})$ plane, stretching from the point $C^{\prime}$ to the point where it intersects with the curve of the extremal capped black hole limit, illustrated as $F$ in Fig. 4(c), where
@@
    Figure 5: $q_{{\mathcal{D}}}$ on the $j_{\phi}=0.2404$ phase near the extremal capped black hole limit (black dot). The local maximum (blue dot) corresponds to the blue dot-dashed curve in Fig. 4(c).

### v2

@@ IV Physical properties
Finally, we study the phase space of the new solution. The new capped black hole solution is determined by four independent parameters $(\ell,\sigma,\tilde{\nu},\tilde{\gamma})$. Here, we focus on the phase diagram for the $n=1$ case with $\sigma=0.8$. Other cases with $\sigma\not=0.8$ are qualitatively similar.
@@ IV.1 Thermodynamics
$\Phi_{\mathcal{D}}$ $:=-i_{v_{\rm in}}A\biggr{|}_{x=1}=-(A_{\phi}+nA_{\psi})\biggr{|}_{x=1}=-\frac{\sqrt{3}\ell cs\tilde{v}_{0}\left(2c(1-\gamma)f_{2}\nu\tilde{g}_{1}(a-b)-f_{1}s\left(c_{3}f_{4}f_{17}-bc_{2}f_{5}f_{13}\right)\right)}{2\nu(1-\gamma)f_{1}f_{2}},$ (96)
@@
with $\Phi(y):=-i_{v_{\cal H}}A|_{x=1}=-(A_{t}+\omega_{\psi}A_{\psi}+\omega_{\phi}A_{\phi})|_{x=1}$.
Figure 4: The $\sigma=0.8$ phase in the $(j_{\psi},j_{\phi})$, $(\phi_{\mathcal{D}},j_{\phi})$ and $(q_{\mathcal{D}},j_{\phi})$ planes. The $\tilde{\nu}={\rm constant}$ phases are shown by thin blue curves. To keep the visibility in the panel (a), the $\tilde{\nu}={\rm constant}$ curves are displayed only in the close-up. In each panel, the limits to the extremal capped black hole are depicted with black and blue curves. The limits to the extremal black hole without a bubble are represented by orange and green curves. Additionally, the critical curves for the large and small bubble branches are indicated by red dot-dashed lines, which also form part of the boundary in panel (a). In panel (a), two branches exist for given $(j_{\psi},j_{\phi})$ within the purple region, while only the large bubble branch is present in the gray region. In panels (b) and (c), the critical curves separate the two branches, with the orange and green dotted curves marking the transition between regions with single and double branches. The red regions denote the small bubble branch and the blue regions correspond to the overlapping large bubble branch from the purple region in panel (a). In both panels (b) and (c), only the large bubble branch exists within the gray areas. In each panel, the two limit curves to the extremal capped black hole and extremal bubbleless black hole intersect at points $B$ and $B^{\prime}$. The critical curve terminates at points $D$ and $D^{\prime}$ on the extremal bubbleless black hole limits. In panels (a) and (b), the two limits to extremal capped black hole join at the point $C=C^{\prime}$, while, in panel (c), two limits terminate at different points $C$ and $C^{\prime}$ linked by the $j_{\phi}=0.240\ldots$ line (gray dashed line). Additionally, the phase diagram in panel (c) is delineated by the $\partial_{\tilde{\nu}}j_{\phi}\partial_{\tilde{\gamma}}q_{\mathcal{D}}-\partial_{\tilde{\nu}}q_{\mathcal{D}}\partial_{\tilde{\gamma}}j_{\phi}=0$ curve (blue dot-dashed curve), which begins at $F$ on the lower extremal capped black hole limit and ends at $E$ on the $j_{\phi}=0.240\ldots$ line.
@@
In Fig. 4, the phase for $n=1$ and $\sigma=0.8$ is depicted in the $(j_{\psi},j_{\phi})$, $(\phi_{\mathcal{D}},j_{\phi})$, and $(q_{\mathcal{D}},j_{\phi})$ planes, where the solution exists within the colored regions. Since the BPS bound is saturated only at the limit $\sigma\to\pm 1$, all the phases for $\sigma=0.8$ are non-BPS (non-supersymmetric). In each panel, curves of constant $\tilde{\nu}$ are shown as thin blue lines. The capped black hole solutions are not uniquely determined by their conserved charges $(M,Q,J_{\psi},J_{\phi})$ and feature two distinct branches, which we call large bubble branch and small bubble branch. As explained in detail below, the former always has the larger area of a bubble than the latter. In Fig. 4(a), both branches coexist within the purple region. These branches are distinguishable by their magnetic flux or potential values. In Figs. 4(b) and (c), the small bubble branch is represented by the red region, while the large bubble branch occupies both blue and gray regions,
The solution of Eqs. (78) and (81) in the range (90) provides only a part of the phase diagram if all parameters remain real (gray region in Fig. 5). However, it turns out that the solution for $n=-1$ and $\sigma=-0.8$ obtained from Eqs. (88) and (89) in the same range (90) compensates the lacking region via the transformation (92) (blue hatched region in Fig. 5). These two regions overlap near the $\tilde{\gamma}=0$ curves of each region (black and blue dashed curves in Fig. 5). Both solutions coincide in the overlap region, and hence the entire phase diagram is consistently covered by two regions.
Figure 5: The extension of the $n=1,\sigma=0.8$ phase by using the $n=-1,\sigma=-0.8$ phase. The red region represents the $n=1,\sigma=0.8$ phase within the parameter bounds $1\leq\tilde{\nu}\leq\sqrt{2}$ and $0\leq\tilde{\gamma}\leq\sqrt{2-\tilde{\nu}^{2}}$. The blue hatched region illustrates the phases derived from the $n=-1,\sigma=-0.8$ phase, for the same ranges of $\tilde{\nu}$ and $\tilde{\gamma}$, obtained by the transformation $(t,\psi)\to(-t,-\psi)$. The upper boundary of the former region (indicated by a red dashed curve) and the lower boundary of the latter (represented by a blue dashed curve) are defined by the limit $\tilde{\gamma}\to 0$ in each respective solution. Notably, there is a slight overlap between these two regions within these boundaries.
@@
  • The limit of non-BPS (nonsupersymetric) extremal black hole without bubbles (orange and green solid curves in Fig. 4), as $\tilde{\gamma}\to\sqrt{2-\tilde{\nu}^{2}}$ with $\tilde{\nu}$ held constant, where the surface gravity $\kappa$ approaches zero. In this limit, the angular momenta correspond to those of the extremal Cvetič-Youm black holes, maintaining the same charge-to-mass ratio as specified in Eq. (82). For the limit with smaller $j_{\phi}$ (orange curves in Fig. 4), the angular momenta are given by
@@
    $=\frac{\left(\sigma^{3}+1\right)\left(3\left(\sigma^{3}-3\right)\tilde{\nu}^{2}(\tilde{\nu}^{2}-2)+2\left(\sqrt{\sigma^{6}\left(-3\tilde{\nu}^{4}+6\tilde{\nu}^{2}+1\right)+2\sigma^{3}\left(3\tilde{\nu}^{4}-6\tilde{\nu}^{2}+1\right)+1}-\sigma^{3}-1\right)\right)}{9\tilde{\nu}^{2}\left(2-\tilde{\nu}^{2}\right)},$ (100)
    and the entropy is given by
@@
    ×$\quad\times\sqrt{\sigma^{6}\left(-3\tilde{\nu}^{4}+6\tilde{\nu}^{2}+1\right)+2\sigma^{3}\left(3\tilde{\nu}^{4}-6\tilde{\nu}^{2}+1\right)+1}\biggr{]}^{1/2}.$ (101)
    For the limit with larger $j_{\phi}$ (green curves in Fig. 4), the angular momenta are given by
    $j_{\psi}^{\rm BH}=\frac{j_{1}-\sigma^{3}j_{2}}{(1+\sigma^{2})^{3/2}},\quad j_{\phi}^{\rm BH}=-\frac{j_{2}-\sigma^{3}j_{1}}{(1+\sigma^{2})^{3/2}},$ (102)
    where
    $j_{1}$ $=1+j_{2}$
    $=-\frac{\left(1-\sigma^{3}\right)\left(\left(1-\sigma^{3}\right)\left(\tilde{\nu}^{4}-2\tilde{\nu}^{2}+2\right)-2\sqrt{\sigma^{6}\left(\tilde{\nu}^{2}-1\right)^{2}-2\sigma^{3}\left(\tilde{\nu}^{2}-1\right)^{2}+1}\right)}{\tilde{\nu}^{2}\left(2-\tilde{\nu}^{2}\right)},$ (103)
    and the entropy is given by
    $a_{\cal H}^{\rm BH}=\frac{2\sqrt{2}\left(1-\sigma^{3}\right)^{3/2}}{\left(\sigma^{2}+1\right)^{3/2}\left(2-\tilde{\nu}^{2}\right)^{2}}\biggr{[}\sigma^{9}\tilde{\nu}^{2}\left(\tilde{\nu}^{2}-1\right)-3\sigma^{6}\tilde{\nu}^{2}\left(\tilde{\nu}^{2}-1\right)+2\sigma^{3}\left(\tilde{\nu}^{4}-\tilde{\nu}^{2}+1\right)-2$
    $\quad+\left(2-\sigma^{3}\left(2-\sigma^{3}\right)\tilde{\nu}^{2}\right)\sqrt{\sigma^{6}\left(\tilde{\nu}^{2}-1\right)^{2}-2\sigma^{3}\left(\tilde{\nu}^{2}-1\right)^{2}+1}\biggr{]}^{1/2}.$ (104)
    $\phi_{\mathcal{D}}$ and $q_{\mathcal{D}}$ take the same form on these curves
@@
    with $(j_{1},j_{2})$ in Eq. (100) or in Eq. (103). Since the Cvetič-Youm black hole does not possess the bubble of $S^{2}$ outside the horizon, both the magnetic flux and magnetic potential are trivially zero. We note that the limit of the entropy (101) does not coincide with that of of the corresponding Cvetič-Youm black hole in Eq. (106). The similar discontinuity is reported for the black hole limit in the black ring phase Elvang:2007hs ; Elvang:2004rt . As speculated in Ref. Elvang:2004rt , this may be because the spacetime topology changes at the limit.
    The two extremal black hole curves intersect at the limit $\tilde{\nu}\to\sqrt{2}$ ($A$ in Fig. 4), where the angular momenta and entropy become
    $(j_{\psi},j_{\phi})\bigr{|}_{A}$ $=\frac{(1,\sigma^{3})}{(1+\sigma^{2})^{3/2}},$ (107)
    $a_{{\cal H}}\bigr{|}_{A}$ $=0.$ (108)
    The common endpoint $A$ corresponds to the extremal Cvetič-Youm phase with zero entropy. Each curve has another endpoint at the limit $\tilde{\nu}\to 1$ ($B$ and $B^{\prime}$ in Fig. 4), where the angular momenta and entropy become
    $(j_{\psi},j_{\phi})\bigr{|}_{B}$ $=\frac{(j_{1}+\sigma^{3}j_{2},j_{2}+\sigma^{3}j_{1})}{(1+\sigma^{2})^{3/2}}\biggr{|}_{B},$ (109)
    $j_{1}\bigr{|}_{B}$ $=1+j_{2}\bigr{|}_{B}=\biggr{\{}\begin{array}[]{cc}1-9^{-1}(2-\sigma^{3})^{2}&(2^{-1/3}<\sigma<1)\\ 1-\sigma^{6}&(-1<\sigma<2^{-1/3})\end{array},$ (112)
    $a_{{\cal H}}\bigr{|}_{B}$ $=2\sqrt{2}\left(\frac{1+\sigma^{3}}{1+\sigma^{2}}\right)^{3/2}\times\quad\biggr{\{}\begin{array}[]{cc}9^{-1}(2-\sigma^{3})\sqrt{2\sigma^{3}-1}&(2^{-1/3}<\sigma<1)\\ \sqrt{\sigma^{6}(1-2\sigma^{3})}&(-1<\sigma<2^{-1/3})\end{array}.$ (115)
@@
    $(j_{\psi},j_{\phi})\bigr{|}_{B^{\prime}}$ $=\frac{(j_{1}-\sigma^{3}j_{2},-j_{2}+\sigma^{3}j_{1})}{(1+\sigma^{2})^{3/2}}\biggr{|}_{B^{\prime}},$ (116)
    $j_{1}\bigr{|}_{B^{\prime}}$ $=1+j_{2}\bigr{|}_{B^{\prime}}=1-\sigma^{6},$ (117)
    $a_{{\cal H}}\bigr{|}_{B^{\prime}}$ $=2\sqrt{2}\,|\sigma|^{3}\left(\frac{1-\sigma^{3}}{1+\sigma^{2}}\right)^{3/2}.$ (118)
  • The limit to a non-BPS (nonsupersymetric), extremal black hole limit with a bubble, i.e., an extremal capped black hole (black and blue curves in Fig. 4) as $\tilde{\nu}\to 1$ and $\tilde{\gamma}\to 1$ with
    $\tilde{\gamma}=1-(\delta/\mu^{2})\beta,\quad\tilde{\nu}=1+\delta\beta,\quad(0<\mu<1).$ (119)
    For the limit with smaller $j_{\phi}$ (with larger $j_{\phi}$), substituting these expressions into Eqs. (78) and (81) (Eqs. (88) and (89)), and subsequently taking the limit $\beta\to 0$, we derive these two equations for the parameters $(\sigma,\tilde{b},\delta,\mu)$, solvable for a given $(\sigma,\mu)$ configuration. These limits intersects with the previous extremal black hole limit at $\mu\to 1$ ($B$ and $B^{\prime}$ in Fig. 4) and terminates at $\mu\to 0$ ($C$ and $C^{\prime}$ in Fig. 4). At the latter points, the spacetime corresponds to the extremal phase with zero entropy. For instance, at $\sigma=0.8$, the physical parameters at $C$ are numerically determined as follows:
@@
    At the point $C^{\prime}$, the values of $j_{\psi},j_{\phi}$ and $\phi_{\mathcal{D}}$ are the same as those at the point $C$, but $q_{{\mathcal{D}}}$ are different:
    $q_{\mathcal{D}}\bigr{|}_{C^{\prime}}=-0.0455....$ (121)
@@
    where $j_{\psi}$ has a local minimum on a $j_{\phi}={\rm constant}$ surface. This curve forms part of the boundary in the phase diagram only in the $(j_{\psi},j_{\phi})$ plane, as shown in Fig. 4(a). In the $(\phi_{\mathcal{D}},j_{\phi})$ and $(q_{\mathcal{D}},j_{\phi})$ planes, it distinguishes between the large and small bubble branches, as illustrated Fig. 4(b) and (c). This curve ends on the extremal black hole curves ($D$ and $D^{\prime}$ in Fig. 4). The respective physical quantities at these end points are numerically determined as
    $(j_{\psi},j_{\phi},\phi_{{\mathcal{D}}},q_{{\mathcal{D}}})\bigr{|}_{D}$ $=(0.420...,0.187...,0.866...,0),$ (123)
    $(j_{\psi},j_{\phi},\phi_{{\mathcal{D}}},q_{{\mathcal{D}}})\bigr{|}_{D^{\prime}}$ $=(0.416...,0.303...,0.866...,0).$ (124)
    The line of $j_{\phi}=0.240\dots$ extending from $C$ to $C^{\prime}$ in the $(q_{\mathcal{D}},j_{\phi})$ plane (gray dashed line in Fig. 4(c)). At $\tilde{\nu}\to 1$, the constant curves of $\tilde{\gamma}$ with different values end at different points on this line. In the $(j_{\psi},j_{\phi})$ and $(\phi_{\mathcal{D}},j_{\phi})$ planes, this curve degenerates to the single point ($C=C^{\prime}$ in Fig. 4(a) and (b)), with the value of $\tilde{\gamma}$ determining the direction in which the $\tilde{\gamma}={\rm constant}$ curve approaches this point.
@@
    where $q_{\mathcal{D}}$ has a local maximum on the $j_{\phi}={\rm constant}$ phase (Fig. 6). As in Fig. 4(c), this curve forms a part of the boundary in the $(q_{\mathcal{D}},j_{\phi})$ plane, stretching from the point $E$ on the line of $j_{\phi}=0.240\dots$ to the point $F$ at which it intersects with the curve of the extremal capped black hole limit, as illustrated in Fig. 4(c), where
    $(j_{\psi},j_{\phi},\phi_{{\mathcal{D}}},q_{{\mathcal{D}}})\bigr{|}_{E}=(0.482...,0.240...,0.780...,-0.0420...),$ (126)
@@
    Figure 6: $q_{{\mathcal{D}}}$ on the $j_{\phi}=0.2404$ phase near the extremal capped black hole limit (black dot) for $n=1$ and $\sigma=0.8$. The local maximum (blue dot) corresponds to the blue dot-dashed curve in Fig. 4(c).